# Efficient Unstructured Pruning of Mamba State-Space Models for Resource-Constrained Environments

### v1

@@ Appendix C Fine-Grained Ablations and Extended Results
This section presents ablation studies and extended results moved from the Results section, including fine-grained component ablations and cross-dataset performance.
C.1 Fine-Grained Ablations
We analyze the impact of gradient exponent, global vs. layer-wise pruning, pruning schedules, and component-specific pruning.
C.1.1 Gradient Exponent Ablation
Table 8 shows the impact of $\alpha$ on WikiText-103 perplexity for Mamba-Base at 50% sparsity.
Table 8: Perplexity on WikiText-103 for different $\alpha$ values at 50% sparsity.
$\alpha$ Perplexity Inference (ms/token)
0.0 (Magnitude Only) 21.5 0.87
0.5 21.0 0.84
1.0 (Default) 20.7 0.82
2.0 20.9 0.83
$\alpha=1.0$ balances magnitude and gradient, minimizing perplexity.
C.1.2 Global vs. Layer-Wise Pruning
Table 9 compares global and layer-wise pruning.
Table 9: Perplexity on WikiText-103 for global vs. layer-wise pruning at 50% sparsity.
Strategy Perplexity Params Saved (%)
Layer-Wise 21.2 50
Global 20.7 50
Global pruning outperforms by 0.5 points, optimizing parameter allocation.
C.1.3 Pruning Schedule
Table 10 evaluates schedule functions.
Table 10: Perplexity on WikiText-103 for different pruning schedules at 50% sparsity.
Schedule Perplexity Training Time (h)
Linear 21.4 28
Exponential 21.1 30
Cubic (Ours) 20.7 29
The cubic schedule minimizes perplexity with comparable training time.
C.1.4 Fine-Grained Component Ablation
Table 11 details pruning specific Mamba components at 70% sparsity within the component.
Table 11: Perplexity on WikiText-103 when pruning individual components at 70% sparsity within the component.
Component Type Parameter Perplexity Perplexity
Percentage Increase (%)
Dense Model (Reference) 100% 19.8 0.0%
SSM Components
$\Delta$ (Time-step) Projections 4% 20.3 2.5%
$A_{\text{log}}$ Projections 4% 21.6 9.1%
$B$ Projections 4% 20.5 3.5%
$C$ Projections (Output) 4% 20.1 1.5%
Linear Components
Input Projections 12% 21.2 7.1%
Output Projections 12% 21.0 6.1%
Gating Networks 8% 21.4 8.1%
Other Components
Layer Norm Parameters 1% 19.9 0.5%
Embedding Layer 30% 22.5 13.6%
Output Head 21% 21.8 10.1%
Figure 9: Optimized non-uniform pruning allocation. (A) Component-specific sparsity levels at 70% global sparsity. (B) Performance comparison between uniform and optimized allocation. (C) Parameter distribution before and after pruning.
C.2 Extended Results
@@
C.2.1 Cross-Dataset Performance
@@
Table 12: Performance metrics for Mamba-Base at 50% sparsity across datasets.
@@
C.2.2 Supplementary Figures
@@
Figure 11: Robustness to perturbations across sparsity levels for Mamba-Base.
Appendix D Related Work and Future Work
This section expands the Related Work and details future research directions.
D.1 Related Work
Early sequence models used RNNs [52], with LSTMs [52] and GRUs [53] addressing vanishing gradients. Transformers [1] surpassed RNNs but face quadratic complexity [4]. Alternatives like RetNet [51], RWKV [50], and hybrid models [38] balance efficiency and performance. SSMs, including S4 [6], DSS [7], and Mamba [8], offer linear-time complexity, with Mamba’s selective mechanism excelling on dense data [9, 12].
Pruning has been applied to CNNs [17] and Transformers [19], but SSM-specific pruning is underexplored due to stability needs [22]. Our work addresses this gap.
D.2 Future Work
Several promising research directions emerge from our findings. Combining pruning with quantization techniques [36] could yield multiplicative efficiency gains, potentially enabling deployment on even more constrained devices. Knowledge distillation approaches [37] offer opportunities to use dense Mamba models to train even more effective sparse ones. Exploring pruned hybrid SSM-attention architectures [38] could yield optimal balances of efficiency and expressiveness. We also see value in developing adaptive hyperparameter optimization (e.g., for the $\alpha$ parameter) to further streamline the pruning process. From a theoretical perspective, deeper exploration of the connections between prunability and generalization [27] could yield fundamental insights into sparse SSMs. Finally, hardware-specific optimizations leveraging sparse tensor accelerators [35] present opportunities for additional real-world performance improvements.
Appendix E Additional Supplementary Material
This section provides extra figures, implementation details, and extended metrics.
E.1 Additional Figures
Figure 12: Cross-layer importance score distribution for Mamba-Base on WikiText-103, showing higher importance in early and middle layers.
Figure 12 complements Figure 4, highlighting layer-wise importance.
E.2 Pseudo-Code
Algorithm 1 Gradient-Aware Pruning for Mamba Models
1:  Input: Model $\theta$, dataset $\mathcal{D}$, target sparsity $s_{f}$, pruning steps $T$, gradient exponent $\alpha$, learning rate $\eta$
2:  Initialize mask $M\leftarrow\mathbf{1}$, sparsity $s_{0}\leftarrow 0$
3:  for $t=t_{0}$ to $T$ do
4:     Compute gradients $\frac{\partial\mathcal{L}}{\partial\theta}$ on $\mathcal{D}$
5:     Compute importance scores $S(\theta_{ij})=|\theta_{ij}|\cdot\left|\frac{\partial\mathcal{L}}{\partial\theta_{ij}}\right|^{\alpha}$
6:     Update sparsity $s_{t}=s_{f}+(s_{0}-s_{f})\left(1-\frac{t-t_{0}}{T-t_{0}}\right)^{3}$
7:     Sort $S(\theta_{ij})$ globally, mask lowest $s_{t}$-fraction to 0 in $M$
8:     Apply mask: $\theta\leftarrow\theta\odot M$
9:     Fine-tune $\theta$ with AdamW ($\eta$) for 5K steps
10:     Compute stability score $S_{\text{stab}}(M)=\max_{i}|\lambda_{i}(\bar{A}\odot M)|-1$
11:     if $S_{\text{stab}}(M)>\epsilon$ then
12:        Adjust mask to prioritize SSM parameters
13:     end if
14:  end for
15:  Output: Pruned model $\theta$, mask $M$
E.3 Extended Robustness Metrics
Table 13 extends Table 6 with additional text classification robustness evaluation metrics.
Table 13: Text classification accuracy (%) under perturbations for Mamba-Base.
Model Clean Synonym Swap Char Swap Word Deletion Avg Drop
Dense 93.2 73.8 70.2 76.8 19.8
Ours-Pruned (30%) 92.8 75.1 72.5 78.2 17.4
Ours-Pruned (50%) 92.6 74.2 71.8 77.8 16.6
Ours-Pruned (70%) 91.5 72.9 70.5 76.4 18.2
Table 14: Extended language modeling robustness evaluation on WikiText-103, expanding on Table 7.
Model Adversarial Spelling Errors Paraphrase Mix
Mamba (Dense) 76.8% 83.5% 88.1% 74.2%
Mamba (Magnitude, 50%) 73.2% 81.7% 85.8% 71.6%
Mamba (Ours, 50%) 79.4% 85.3% 90.2% 76.8%
These results demonstrate that our pruned models consistently outperform both dense and magnitude-pruned models across all robustness metrics, with the advantage particularly pronounced for adversarial perturbations (2.6% improvement over dense) and combined perturbations (2.6% improvement over dense).
E.4 Extended Efficiency Metrics
Table 15 extends Table 5.
Table 15: Memory usage (GB) and latency (ms/token) for Mamba-Base at 50% sparsity.
Platform Dense Memory Pruned Memory Dense Latency Pruned Latency
NVIDIA A100 7.2 3.9 1.45 0.82
NVIDIA V100 8.1 4.3 1.62 0.91
Edge (Jetson TX2) 6.8 3.7 12.5 7.1
Appendix F Additional Analysis and Implementation Details
F.1 Stability Threshold and Sensitivity
@@
The stability threshold $\epsilon$ in our stability score calculation (Section 3.4) serves as a safety margin to ensure eigenvalues remain within the unit circle. Through empirical testing across different Mamba models, we find that values in the range $\epsilon\in[0.005,0.02]$ work well, with $\epsilon=0.01$ providing a good balance between stability enforcement and pruning flexibility.
@@
Figure 13: Impact of stability threshold $\epsilon$ on model performance. (A) Perplexity vs. $\epsilon$ value at 50% sparsity, showing optimal performance in the range [0.005, 0.02]. (B) Percentage of affected parameters requiring stability correction at different $\epsilon$ values. (C) Eigenvalue distribution for different $\epsilon$ settings, demonstrating how stricter thresholds constrain the eigenvalue magnitudes.
F.2 Reducing Computational Overhead
@@
Rather than computing gradients for every iteration, we found that accumulating gradients over 5-10 batches before updating importance scores yields similar results while reducing computational overhead by 3-5x for this component. Instead of pruning at every step of the schedule, applying pruning every k iterations (where k scales with batch size) maintains comparable performance while significantly reducing training time. For our experiments, pruning every 50-100 iterations worked well for large batch sizes. We also experimented with more aggressive cubic schedules that complete pruning in 60
With these optimizations, we reduced the training overhead from the reported 2.5x to approximately 1.7x while maintaining performance within 0.5
F.3 Component Sensitivity Analysis
@@
F.4 Adaptive Gating and Pruning
The cross-architecture experiments (Appendix B.4, Figure B.3) demonstrate that models with adaptive gating mechanisms (Mamba and GSS) show better pruning tolerance than fixed-dynamics SSMs. Our analysis confirms that the adaptive gating parameters themselves are indeed critical for maintaining performance under pruning.

### v2

@@ Appendix C Fine-Grained Ablations and Extended Results
The results of our ablation studies are summarized in Table 8 in the main text (Section 5.2). Our ablation experiments confirm that global pruning outperforms layer-wise pruning, a balanced gradient-magnitude importance score ($\alpha=1.0$) is superior to pure magnitude-based pruning, and a cubic schedule for increasing sparsity yields better performance than linear or exponential schedules.
Figure 9: Comparison of different pruning schedules showing sparsity progression, performance impact, and parameter adaptation dynamics.
Figure 10: Component-wise sensitivity analysis showing performance impact and recovery dynamics for different component types.
Figure 11: Optimized non-uniform pruning allocation showing component-specific sparsity levels and performance comparisons.
C.1 Extended Results
@@
C.1.1 Cross-Dataset Performance
@@
Table 11: Performance metrics for Mamba-Base at 50% sparsity across diverse datasets and tasks.
@@
Figure 12: Time-series forecasting results showing MSE across horizons and performance retention at different sparsity levels.
The cross-dataset evaluation reveals consistent patterns in the effectiveness of our pruning approach across diverse domains. For language modeling tasks, our method consistently achieves 50% parameter reduction with only a 4-5% relative decrease in performance across all datasets, compared to 8-10% degradation with standard magnitude pruning. This consistency across datasets of varying complexity (from The Pile to PG-19) demonstrates the robustness of our approach to different linguistic distributions and contexts.
Long-range sequence tasks show particularly strong results, with our pruning method maintaining performance within 0.7 percentage points of dense models on Path-X, compared to 1.7 points for magnitude pruning. This suggests our approach better preserves the selective state-space mechanisms critical for modeling dependencies in long sequences. The ListOps task, which requires hierarchical reasoning, shows slightly larger degradation (0.7%) but still outperforms magnitude pruning by 1.4 percentage points.
For time-series forecasting, our pruned models achieve MSE values within 2.7% of dense models, compared to 5.2% degradation for magnitude pruning. This trend holds across different forecasting horizons, with performance gaps widening for longer-range predictions (96h vs. 48h), highlighting our method’s advantage in preserving long-range dependencies.
Audio processing tasks show remarkable resilience to pruning, with our method maintaining performance within 0.4-0.8 percentage points across all datasets. Speech Commands, which involves simple classification, shows the smallest degradation (0.4%), while LibriSpeech, which requires more complex sequence modeling, shows slightly larger impacts (0.3% WER increase). These results suggest that audio tasks may be particularly amenable to pruning due to inherent redundancy in audio representations.
Vision tasks exhibit slightly larger performance drops compared to other domains (0.6-0.8 percentage points), but still significantly outperform magnitude pruning. The more complex the task, the larger the gap between our method and magnitude pruning—for instance, in semantic segmentation (ADE20K), our approach outperforms magnitude pruning by 1.8 percentage points, compared to 1.2 points for classification tasks. This suggests that tasks requiring fine-grained spatial understanding benefit more from our gradient-aware approach.
Across all domains, we observe that the performance gap between our method and magnitude pruning widens as task complexity increases, indicating that gradient information becomes increasingly valuable for identifying critical parameters in more challenging contexts.
C.1.2 Supplementary Figures
@@
Figure 13 reveals the non-uniform distribution of parameters retained after pruning across different layers and component types. Several important patterns emerge from this analysis. First, we observe that earlier layers (layers 1-3) retain approximately 10-15% more parameters than later layers, consistent with findings in other architectures where early feature extraction requires more capacity. Second, within each layer, SSM parameters show significantly higher retention rates (65-75%) compared to linear projection parameters (30-40%), confirming our hypothesis that state-space components are more critical to model performance.
A notable finding is the differential impact across SSM component types: state transition parameters ($A$ matrices) show the highest retention rates (70-80%), followed by input projection parameters ($B$ matrices, 60-70%), and finally output projection parameters ($C$ matrices, 50-60%). This ordering aligns with theoretical understanding of SSMs, where state transitions most directly influence the model’s ability to capture temporal dependencies. The selective mechanism parameters ($\Delta$ projections) also show high retention rates (65-75%), demonstrating their importance for Mamba’s data-dependent processing.
The parameter distribution also reveals interesting patterns across attention heads within each layer. We find that different heads specialize in capturing different dependency ranges, with some heads showing near-zero pruning while others are pruned more aggressively. This supports the hypothesis that Mamba’s selective attention mechanism creates specialization across different components, with our gradient-aware pruning preserving this functional diversity.
@@
Figure 14: Robustness to perturbations across sparsity levels for various input perturbation types.
Figure 14 extends our robustness analysis by examining how different levels of sparsity affect Mamba’s resilience to various input perturbations. The figure shows accuracy retention (as a percentage of clean performance) across five perturbation types at sparsity levels ranging from 0% (dense) to 80%.
Our pruned models demonstrate superior robustness to the dense baseline at moderate sparsity levels (30-50%), with a peak improvement of 2.8% at 50% sparsity. This unexpected enhancement can be attributed to the regularizing effect of pruning, which reduces the model’s tendency to overfit to specific input patterns. The improvement is most pronounced for synonym substitutions and word insertions (3.5-4.0% gain), suggesting that pruning helps the model focus on semantic understanding rather than exact token matching.
As sparsity increases beyond 60%, robustness advantages diminish and eventually reverse, with dramatic drops at extreme sparsity (>75%). This indicates a "sweet spot" around 50% sparsity where the regularization benefits of pruning outweigh the capacity reduction. The robustness curves also reveal that different perturbation types exhibit different sensitivity patterns: character-level perturbations show earlier degradation as sparsity increases compared to word-level perturbations, suggesting that character-level processing requires more model capacity.
Interestingly, models pruned with our gradient-aware method maintain significantly better robustness compared to magnitude-pruned models at all sparsity levels, with gaps widening at higher sparsity. This confirms that our pruning approach better preserves the model’s generalization capabilities by retaining parameters critical for capturing underlying patterns rather than memorizing specific inputs.
Appendix D Additional Analysis and Implementation Details
D.1 Stability Threshold and Sensitivity
@@
The stability threshold $\epsilon$ in our stability score calculation (Section 3.1.4) serves as a safety margin to ensure eigenvalues remain within the unit circle. Through empirical testing across different Mamba models, we find that values in the range $\epsilon\in[0.005,0.02]$ work well, with $\epsilon=0.01$ providing a good balance between stability enforcement and pruning flexibility.
@@
Figure 15: Impact of stability threshold $\epsilon$ on model performance, parameter correction rates, and eigenvalue distributions.
D.2 Reducing Computational Overhead
@@
Rather than computing gradients for every iteration, we found that accumulating gradients over 5-10 batches before updating importance scores yields similar results while reducing computational overhead by 3-5x for this component. Instead of pruning at every step of the schedule, applying pruning every k iterations (where k scales with batch size) maintains comparable performance while significantly reducing training time. For our experiments, pruning every 50-100 iterations worked well for large batch sizes. We also experimented with more aggressive cubic schedules that complete pruning in 60% of training rather than 75%, finding a modest 0.5-1.0% performance drop but a 20% reduction in additional training time.
With these optimizations, we reduced the training overhead from the reported 2.5x to approximately 1.7x while maintaining performance within 0.5% of our primary results. For deployment scenarios where training efficiency is paramount, these trade-offs provide practical alternatives.
D.3 Component Sensitivity Analysis
@@
Figure 16: Distribution of parameter importance scores across model components and their evolution during training.
@@
D.4 Adaptive Gating and Pruning
The cross-architecture experiments (Appendix D.7) demonstrate that models with adaptive gating mechanisms (Mamba and GSS) show better pruning tolerance than fixed-dynamics SSMs. Our analysis confirms that the adaptive gating parameters themselves are indeed critical for maintaining performance under pruning.